OpenStudy (anonymous):

Write an equation in slope-intercept form for the line with slope 1/2 that contains the poiint (4,6)

OpenStudy (anonymous):

y-6=1/2(x-4) y = 1/2x -2+6 y=1/2x+4

OpenStudy (anonymous):

Thank you! :)